OpenStudy (anonymous):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 25 feet and the child ride has a radius of 10 feet, what is the ratio between the volumes of each ride?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@jhonyy9

OpenStudy (amistre64):

like a cone adults similar, smaller version for children. If the adult ride has a radius of 25 feet and the child ride has a radius of 10 feet, what is the ratio between the volumes of each ride? Can you define the formula for the volume of a cone?

OpenStudy (anonymous):

no I don't know the formal for volume of a cone could u tell me

OpenStudy (amistre64):

i could; but it would be better if you could find it in your material, or at least google it .. that way you have more weighted participation in the process.

OpenStudy (anonymous):

surface area

OpenStudy (anonymous):

I mean if u tell all I got to do is pug it in right

OpenStudy (amistre64):

if we know the formula, then the rest is just childs play :)

OpenStudy (anonymous):

v=1/3b*h

OpenStudy (amistre64):

thats a good formula, but lets be sure about it: b = base area

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

in this case, the adults have a volume of:\[V_a=\frac13\pi~r^2h\] its similar to some scaled version of:\[V_c=\frac13\pi~(rk)^2~(hk)\] which is k^3:1 in ratio

OpenStudy (amistre64):

the key now is to determine the scaling factor: 25k = 10

OpenStudy (anonymous):

what do I do with 25k=10

OpenStudy (anonymous):

solve that

OpenStudy (amistre64):

well, since our ratio is k^3:1, we need a value for k ... 25 is scaled to 10 by some factor k ... 25k = 10, solve for k yes

OpenStudy (amistre64):

i spose we could just as easily scaled it from 10 to 25 but the problem alludes to an adult:child ratio instead of the other way around

OpenStudy (anonymous):

I got 2.5

OpenStudy (amistre64):

\[k=\frac 1{2.5}~:~or~k=2.5\] \[k^3=\frac 1{2.5^3}~:~or~k^3=(2.5)^3\] depending on the how we define the setup i spose

OpenStudy (anonymous):

which one would u pick

OpenStudy (amistre64):

hmm, id pick the right side for simplicity: 15.625 to 1 do we have any options to narrow the feild?

OpenStudy (amistre64):

the adult volume is 15.625 times the volume of the childs volume

OpenStudy (anonymous):

is that the answer cause my computer froze and I had to cut it off

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

thats the results we get, so yes. It may have to be formated for whatever is grading it tho

OpenStudy (anonymous):

25:4 15:6 125:8 5:2

OpenStudy (anonymous):

B

OpenStudy (anonymous):

right

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

not B this is asking about the ratio of volumes right?

OpenStudy (anonymous):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 25 feet and the child ride has a radius of 10 feet, what is the ratio between the volumes of each ride? 25:4 15:6 125:8 5:2

OpenStudy (anonymous):

that's whats it asking

OpenStudy (amistre64):

\[\frac13\pi~r^2h\to\frac13\pi~(rk)^2~(hk)\] \[r^2h\to(rk)^2~(hk)\] \[r^2\to(rk)^2~k\] \[1\to k^2~k\] \[1\to k^3\] \[r\to rk\] \[10\to 10k=25;~k=2.5;~k^3=15.625,~15\frac58\] soo\[1: 15+\frac58\] \[8: 8(15)+5\] \[8: 125\]

OpenStudy (amistre64):

125 to 8 then :)

OpenStudy (anonymous):

oh thanks

OpenStudy (amistre64):

youre welcome ... the decimal had me for a loop or a second :)

OpenStudy (anonymous):

lol